mathslover (mathslover):

What is \[S_n - S_{n-1}\]

OpenStudy (anonymous):

\[S= ut +\frac{ 1 }{ 2 }a t^2\] use this and you are done , put t= n and t= n-1

OpenStudy (anonymous):

there will be two equations

mathslover (mathslover):

right and we get : \[\large{S_n = u + \frac{a(2n-1)}{2}}\]

mathslover (mathslover):

but can we prove that : \[S_n - S_{n-1}=S_n\] ?

OpenStudy (anonymous):

No because \[S _{n}\] is the distance traveled in nth second and S(n-1) is the distance traveled in (n-1) seconds, how could their difference be equal to distance traveled in n seconds ... think :)

OpenStudy (anonymous):

*n seconds

OpenStudy (anonymous):

not nth

mathslover (mathslover):

yes very correct but the formula says S_n = u + a(2n-1)/2

mathslover (mathslover):

Oh! OK, so it is S_nth = u + a(2n-1)/2 ?

OpenStudy (anonymous):